OpenStudy (anonymous):

Find the distance from (-4,-6) to the line defined by y=-2x+6. Express as a radical or a number rounded to the nearest hundredth.

OpenStudy (amistre64):

well, the vector of the line is <-2,1> pick a point on the line, and create a vector to the stated point with 2 vectors the angle between them is\[\alpha=cos^{-1}(\frac{u\cdot v}{|u||v|})\] and then the distance form the point is just r sin(a)